OpenStudy (anonymous):

You randomly select one of the urns: X, Y, or Z. You then randomly draw one ball out of the selected urn and note its color. The urn contain the following colored balls: Urn X: 2 red, 3 white, and 2 blue. Urn Y: 1 red, 1 white, 2blue Urn Z: 1 red, 1 white, 2blue 1). If the ball is blue, what is the probability that it was drawn from Urn X____? 2.) If the ball is either blue or white, what is the probability that it was drawn from urn X____?

OpenStudy (icecube):

P(X/B)=probability of getting blue ball from urn X ,if the ball is blue . P(X/B)= p(X∩B)/P(B) = (probability of getting blue ball from Urn X)/Total probability of getting blue ball)  therefore P(X/B) =(1/3*2/7)/(1/3*2/7+1/3*2/4+1/3*2/4) =2/9